# via Object-Centric Visual Representation Learning

### v4

@@ 3.2.2 Object Decomposition
We employ Slot Attention [locatello2020object], inspired by recent advances in object discovery [scouter2021cvpr, zhou2022slot, actionslot2024], as a clustering mechanism to group semantically meaningful patches from $\mathbf{f}_{t}$ into predefined slots, where each slot corresponds to a distinct object region. Unlike previous studies [qian2023semantics, actionslot2024] that sample from a prior distribution, we initialize $N$ object slots (denoted as $\mathbf{s}$) as learnable tokens. We then decompose the feature map $\mathbf{f}_{t}$ of $t$-th frame into $N$ frame-wise object slot features $\mathbf{s}_{t}=\{\mathbf{s}_{t}^{1},\ldots,\mathbf{s}_{t}^{N}\}$, $\mathbf{s}_{t}\in\mathbb{R}^{N\times D{\mathtt{slot}}}$. This design encourages each slot to capture modular semantics [xu2022groupvit, DBLP:conf/iclr/JiaLH23], enabling object-consistent decompositions through its unique competition mechanism that supports maintaining coherent slot representation across frames.
Formally, following the standard slot attention procedure [locatello2020object], we employ three linear transformation heads to map the object slots $\mathbf{s}$ into Query $\mathbf{q}\in\mathbb{R}^{N\times D_{\mathtt{enc}}}$, while frame-wise feature maps $\mathbf{f}_{t}$ into Key $\mathbf{k}\in\mathbb{R}^{H_{\mathtt{enc}}W_{\mathtt{enc}}\times D_{\mathtt{enc}}}$, and Value $\mathbf{v}\in\mathbb{R}^{H_{\mathtt{enc}}W_{\mathtt{enc}}\times D_{\mathtt{enc}}}$. We iteratively calculate attention score and update slot representations via Gated Recurrent Unit (GRU) [DBLP:conf/emnlp/ChoMGBBSB14]. Mathematically, we formulate each iteration as:
@@
Object Prediction. Since each object slot captures multi-level granularity information, using a lightweight prediction head is sufficient for object class, bounding box, and mask prediction. Thus, we employ three lightweight prediction heads, each consisting of a feed-forward network (FFN) with two linear layers–a functional layer followed by a task-specific layer for each output. The classification head outputs object classes $\hat{\mathbf{O}}_{t}=\text{FFN}_{\texttt{cls}}(\mathbf{s}_{t})$, while the box head predicts object coordinates $\hat{\mathbf{B}}_{t}=\text{FFN}_{\texttt{box}}(\mathbf{s}_{t})$. For mask prediction, a lightweight decoder with four transpose convolutions [DBLP:conf/cvpr/LongSD15] upsamples the feature map $\mathbf{f}_{t}$ to the original frame size, resulting in $\mathbf{f^{\prime}}_{t}=\text{Dec}(\mathbf{f}_{t})$. The panoptic mask is then obtained by applying a matrix multiplication between the object slots $\mathbf{s}_{t}$ and the upsampled feature map $\mathbf{f^{\prime}}_{t}$ in the mask head, resulting in a binary mask of each object slot $\hat{\mathbf{m}}_{t}=\text{FFN}_{\texttt{mask}}(\mathbf{s}_{t}\cdot\mathbf{f^{\prime}}_{t})$.
Object Temporal Consistency Learning. Maintaining consistent spatio-temporal representations for object slots is crucial for tracking coherent object features over time. Slot Attention provides a strong basis for object-centric video understanding. However, prior studies have noted that slot-based representations struggle to maintain temporal consistency across consecutive frames in a video [DBLP:conf/nips/ZadaianchukSM23]. To address this limitation, and inspired by prior works [DBLP:conf/cvpr/LiZPCCTL22, DBLP:journals/pami/FischerHPQCDY23], we incorporate object temporal consistency learning using a contrastive loss. Specifically, slots matching the same ground truth index across frames are treated as positive samples, while all others are negatives. Formally, given the $i$-th object slot $\mathbf{s}^{i}_{t}$ at frame $t$, we define $\mathbf{s}^{i}_{t-1}$ as the positive target and $\mathbf{s}^{j}_{t-1}$ as negatives ($j\neq i$). The right side of Fig. 2 illustrates the concept of object temporal consistency learning. The corresponding loss function is formulated as:
@@ 3.2.3 Relation Decomposition
Existing VidSGG approaches [Wang_2024_CVPR] predict relations sequentially after object prediction, leading to high computational complexity and potential inaccuracy if objects are missed. In contrast, UNO predicts objects and relations simultaneously, enabling a parallelized execution that significantly reduces complexity. Using Eq. 4, we apply an enhanced slot attention mechanism to decompose the feature map $\mathbf{f}_{t}$ into relation slots, where each slot captures potential interaction regions. This mechanism enables relation slots to capture the entire spatial context of a frame, rather than being restricted to object pair intersections. This broader coverage significantly enhances relation prediction performance and supports our dynamic triplet mechanism (see Sec. 3.2.4). Specifically, we initialize $K$ relation slots (denoted as $\mathbf{z}$) as learnable tokens. Then, for the $t$-th frame, we obtain frame-wise relation slot features $\mathbf{z}_{t}=\{\mathbf{z}_{t}^{1},\ldots,\mathbf{z}_{t}^{K}\}$, where $\mathbf{z}_{t}\in\mathbb{R}^{K\times D{\mathtt{slot}}}$. Each frame-wise relation slot $\mathbf{z}_{t}$ is then passed through a classification FFN head to predict relations: $\hat{\mathbf{R}}_{t}=\text{FFN}_{\texttt{rel}}(\mathbf{z}_{t})$. Consequently, both relation slots $\mathbf{z}_{t}$ together with its relation class $\hat{\mathbf{R}}_{t}$ and object slots $\mathbf{s}_{t}$ are jointly extracted from $\mathbf{f}_{t}$ at time step $t$.
@@ 3.2.4 Dynamic Triplet Prediction
Existing methods either learn an adjacency matrix [panopticsg2024tpami] to identify subject-object pairs or directly predict pairwise subject-object embeddings [Wang_2024_CVPR], followed by sequential relation prediction. In contrast, our approach introduces a dynamic triplet prediction mechanism that directly associates $N$ objects with $K$ relations, eliminating the need to construct an $N\times N$ object pair matrix. In theory, $K$ can be as large as $N^{2}$, since each object may interact with every other object. However, in practice, $K\ll N$ due to the inherent sparsity of real-world interactions—only a small subset of object pairs exhibit meaningful relationships. This reduces redundancy and duplication in triplet prediction while maintaining high performance. Fig. 3 illustrates the proposed module.
@@
Triplet Duplication Reduction. Unlike prior work [Wang_2024_CVPR], which predicts triplets as a set and often results in duplicate detections, our method leverages slot attention with a built-in competitive mechanism to mitigate redundancy. This design ensures distinct object slots and relation slots focus on appropriate regions, producing unique object and relation predictions for triplets without requiring post-processing steps like Non-Max Suppression (NMS) [DBLP:conf/cvpr/sangBS17].
@@ 3.3 Training Objectives
During training, we first perform Hungarian matching [DBLP:journals/jacm/EdmondsK72] between the predicted and ground-truth object boxes/masks to assign object slots, followed by supervision for detection and classification. The loss function for the DSGG task is defined as:
@@
where $\mathcal{L}_{\text{obj\_cls}}$ is the Cross Entropy (CE) loss, $\mathcal{L}_{\text{box}}$ is the $\ell_{1}$ loss, and $\mathcal{L}_{\text{GIoU}}$ is the GIoU loss [DBLP:conf/cvpr/RezatofighiTGS019].
@@
where $\mathcal{L}_{\text{mask}}$ is the CE loss, and $\mathcal{L}_{\text{dice}}$ is the Dice loss [DBLP:conf/miccai/SudreLVOC17, DBLP:conf/eccv/WangKSJL20].
@@
STTran [cong2021spatial] ICCV’21 ResNet-101 25.2 34.1 37.0 68.6 71.8 71.8 24.6 36.2 48.8 77.9 94.2 99.1
APT [li2022dynamic] CVPR’22 ResNet-101 26.3 36.1 38.3 69.4 73.8 73.8 25.7 37.9 50.1 78.5 95.1 99.2
STTran-TPI [wang2022dynamic] ACM MM’22 ResNet-101 26.2 34.6 37.4 69.7 72.6 72.6 - - - - - -
TR2 [wang2023cross] ICRA’23 ResNet-101 26.8 35.5 38.3 70.9 73.8 73.8 27.8 39.2 50.0 83.1 96.6 99.9
VsCGG [DBLP:conf/mm/LuCSLWH23] ACM MM’23 ResNet-101 27.4 35.8 38.2 70.1 73.4 73.5 29.3 40.2 48.9 78.8 94.9 99.2
TEMPURA [nag2023unbiased] CVPR’23 ResNet-101 28.1 33.4 34.9 68.8 71.5 71.5 29.8 38.1 46.4 80.4 94.2 99.4
DSG-DETR [feng2023exploiting] WACV’23 ResNet-101 30.3 34.8 36.1 - - - 32.1 40.9 48.3 - - -
TPT [zhang2023end] TMM’23 ResNet-101 - - - - - - 32.0 39.6 51.5 85.6 97.4 99.9
@@
OED [Wang_2024_CVPR] CVPR’24 ResNet-50 33.5 40.9 48.9 73.0 76.1 76.1 35.3 44.0 51.8 83.3 95.3 99.2
@@
Table 3: Performance comparison with SOTA PVSG methods on PVSG dataset. The best results are in bold. Next, $\diamondsuit$ and $\heartsuit$ stands for the relation predictor [yang2023pvsg]: 1D Convolution and Transformer Encoder, respectively.
@@
Image Panoptic Segmentation + Tracking [cheng2021mask2former, wang2021different]
@@
Video Panoptic Segmentation [cheng2021mask2former, li2022videoknet]
@@ 4.1 Experimental Settings
Datasets & Evaluation Metrics. We conduct experiments on the Action Genome [ji2020action] for the DSGG task and the PVSG [yang2023pvsg] for the PVSG task. We evaluate the DSGG task following the setting from [Wang_2024_CVPR] and the PVSG task following the setting from [yang2023pvsg]. We train and evaluate the model strictly separately for each benchmark/task, without any data mixing or augmentation strategies.
Implementation Details. We adopt both Vision Transformer (ViT) and Convolutional Neural Networks (CNNs) as the backbone for the Visual Encoder to perform frame-wise feature extraction. More specifically, for ViT, we use ViT-S/14, ViT-B/14, ViT-L/14 [dosovitskiy2020vit] with pre-trained weights from DINO [oquab2023dinov2]; for CNNs, we use ResNet-50 [DBLP:conf/cvpr/HeZRS16] pre-trained weights from MoCo [DBLP:conf/cvpr/He0WXG20]. The number of slots are empirically chosen and can be observed that the optimal number of slots is influenced by both the number of object classes in the dataset and the objects present in the video. Therefore, for DSGG, we set $N=40$ object slots and $K=24$ relation slots; for PVSG, we use $N=96$ object slots and $K=40$ relation slots.
@@ 4.2 Comparison with State of the Arts
Results on Pixel-Level VidSGG (PVSG). The PVSG dataset presents highly dynamic videos and frequent substantial changes in camera angles. Tab. 3 demonstrates that both the Image Panoptic Segmentation + Tracking [cheng2021mask2former, wang2021different] (IPS+T) model and Video Panoptic Segmentation [cheng2021mask2former, li2022videoknet] (VPS) baselines fall short compared to our end-to-end UNO. It is essential to focus on R/mR@20, as it serves as a primary performance metric [yang2023pvsg]. Notably, our model substantially outperforms both IPS+T and VPS at R/mR@20 across mask overlaps of 0.5 and 0.1. For instance, UNO holds the current peak for R@20 at 9.44% in the vIOU threshold of 0.5, indicating that, on average, 01 in every 13 ground-truth triplets is successfully recalled, compared to IPS+T’s best performance of 01 in roughly every 25 triplets (R@20 of 4.51%). However, lowering the threshold to a more lenient 0.1 raises UNO’s score to approximately 17.54%, allowing the model to recall 02 out of every 13 triplets. This suggests that while the model shows higher efficacy than others for recognizing key video content, there remains considerable room for improvement.
@@ 4.3 Ablation Studies
Effect of Multi-Task Learning. Tab. 4 highlights the impact of multi-task learning and cross-task synergy on PVSG, which provides both box-level and pixel-level VidSGG ground truths. AP50 denotes Average Precision@0.5, while PQ is the Panoptic Quality [kirillov2019panoptic]. Using UNO, we observe that training with both bounding boxes and masks improves all metrics, with AP50 increasing from 28.5 to 29.6, PQ from 47.4 to 48.3, and R@20 from 8.03 to 9.44. These results suggest that box-level and pixel-level VidSGG data complement each other, enabling richer representations that enhance individual task performances.

### v5

@@ 3.2.2 Object Decomposition
We employ Slot Attention [33], inspired by recent advances in object discovery [26, 65, 23], as a clustering mechanism to group semantically meaningful patches from $\mathbf{f}_{t}$ into predefined slots, where each slot corresponds to a distinct object region. Unlike previous studies [42, 23] that sample from a prior distribution, we initialize $N$ object slots (denoted as $\mathbf{s}$) as learnable tokens. We then decompose the feature map $\mathbf{f}_{t}$ of $t$-th frame into $N$ frame-wise object slot features $\mathbf{s}_{t}=\{\mathbf{s}_{t}^{1},\ldots,\mathbf{s}_{t}^{N}\}$, $\mathbf{s}_{t}\in\mathbb{R}^{N\times D{\mathtt{slot}}}$. This design encourages each slot to capture modular semantics [58, 20], enabling object-consistent decompositions through its unique competition mechanism that supports maintaining coherent slot representation across frames.
Formally, following the standard slot attention procedure [33], we employ three linear transformation heads to map the object slots $\mathbf{s}$ into Query $\mathbf{q}\in\mathbb{R}^{N\times D_{\mathtt{enc}}}$, while frame-wise feature maps $\mathbf{f}_{t}$ into Key $\mathbf{k}\in\mathbb{R}^{H_{\mathtt{enc}}W_{\mathtt{enc}}\times D_{\mathtt{enc}}}$, and Value $\mathbf{v}\in\mathbb{R}^{H_{\mathtt{enc}}W_{\mathtt{enc}}\times D_{\mathtt{enc}}}$. We iteratively calculate attention score and update slot representations via Gated Recurrent Unit (GRU) [5]. Mathematically, we formulate each iteration as:
@@
Object Prediction. Since each object slot captures multi-level granularity information, using a lightweight prediction head is sufficient for object class, bounding box, and mask prediction. Thus, we employ three lightweight prediction heads, each consisting of a feed-forward network (FFN) with two linear layers–a functional layer followed by a task-specific layer for each output. The classification head outputs object classes $\hat{\mathbf{O}}_{t}=\text{FFN}_{\texttt{cls}}(\mathbf{s}_{t})$, while the box head predicts object coordinates $\hat{\mathbf{B}}_{t}=\text{FFN}_{\texttt{box}}(\mathbf{s}_{t})$. For mask prediction, a lightweight decoder with four transpose convolutions [34] upsamples the feature map $\mathbf{f}_{t}$ to the original frame size, resulting in $\mathbf{f^{\prime}}_{t}=\text{Dec}(\mathbf{f}_{t})$. The panoptic mask is then obtained by applying a matrix multiplication between the object slots $\mathbf{s}_{t}$ and the upsampled feature map $\mathbf{f^{\prime}}_{t}$ in the mask head, resulting in a binary mask of each object slot $\hat{\mathbf{m}}_{t}=\text{FFN}_{\texttt{mask}}(\mathbf{s}_{t}\cdot\mathbf{f^{\prime}}_{t})$.
Object Temporal Consistency Learning. Maintaining consistent spatio-temporal representations for object slots is crucial for tracking coherent object features over time. Slot Attention provides a strong basis for object-centric video understanding. However, prior studies have noted that slot-based representations struggle to maintain temporal consistency across consecutive frames in a video [63]. To address this limitation, and inspired by prior works [28, 14], we incorporate object temporal consistency learning using a contrastive loss. Specifically, slots matching the same ground truth index across frames are treated as positive samples, while all others are negatives. Formally, given the $i$-th object slot $\mathbf{s}^{i}_{t}$ at frame $t$, we define $\mathbf{s}^{i}_{t-1}$ as the positive target and $\mathbf{s}^{j}_{t-1}$ as negatives ($j\neq i$). The right side of Fig. 2 illustrates the concept of object temporal consistency learning. The corresponding loss function is formulated as:
@@ 3.2.3 Relation Decomposition
Existing VidSGG approaches [51] predict relations sequentially after object prediction, leading to high computational complexity and potential inaccuracy if objects are missed. In contrast, UNO predicts objects and relations simultaneously, enabling a parallelized execution that significantly reduces complexity. Using Eq. 4, we apply an enhanced slot attention mechanism to decompose the feature map $\mathbf{f}_{t}$ into relation slots, where each slot captures potential interaction regions. This mechanism enables relation slots to capture the entire spatial context of a frame, rather than being restricted to object pair intersections. This broader coverage significantly enhances relation prediction performance and supports our dynamic triplet mechanism (see Sec. 3.2.4). Specifically, we initialize $K$ relation slots (denoted as $\mathbf{z}$) as learnable tokens. Then, for the $t$-th frame, we obtain frame-wise relation slot features $\mathbf{z}_{t}=\{\mathbf{z}_{t}^{1},\ldots,\mathbf{z}_{t}^{K}\}$, where $\mathbf{z}_{t}\in\mathbb{R}^{K\times D{\mathtt{slot}}}$. Each frame-wise relation slot $\mathbf{z}_{t}$ is then passed through a classification FFN head to predict relations: $\hat{\mathbf{R}}_{t}=\text{FFN}_{\texttt{rel}}(\mathbf{z}_{t})$. Consequently, both relation slots $\mathbf{z}_{t}$ together with its relation class $\hat{\mathbf{R}}_{t}$ and object slots $\mathbf{s}_{t}$ are jointly extracted from $\mathbf{f}_{t}$ at time step $t$.
@@ 3.2.4 Dynamic Triplet Prediction
Existing methods either learn an adjacency matrix [52] to identify subject-object pairs or directly predict pairwise subject-object embeddings [51], followed by sequential relation prediction. In contrast, our approach introduces a dynamic triplet prediction mechanism that directly associates $N$ objects with $K$ relations, eliminating the need to construct an $N\times N$ object pair matrix. In theory, $K$ can be as large as $N^{2}$, since each object may interact with every other object. However, in practice, $K\ll N$ due to the inherent sparsity of real-world interactions—only a small subset of object pairs exhibit meaningful relationships. This reduces redundancy and duplication in triplet prediction while maintaining high performance. Fig. 3 illustrates the proposed module.
@@
Triplet Duplication Reduction. Unlike prior work [51], which predicts triplets as a set and often results in duplicate detections, our method leverages slot attention with a built-in competitive mechanism to mitigate redundancy. This design ensures distinct object slots and relation slots focus on appropriate regions, producing unique object and relation predictions for triplets without requiring post-processing steps like Non-Max Suppression (NMS) [18].
@@ 3.3 Training Objectives
During training, we first perform Hungarian matching [12] between the predicted and ground-truth object boxes/masks to assign object slots, followed by supervision for detection and classification. The loss function for the DSGG task is defined as:
@@
where $\mathcal{L}_{\text{obj\_cls}}$ is the Cross Entropy (CE) loss, $\mathcal{L}_{\text{box}}$ is the $\ell_{1}$ loss, and $\mathcal{L}_{\text{GIoU}}$ is the GIoU loss [44].
@@
where $\mathcal{L}_{\text{mask}}$ is the CE loss, and $\mathcal{L}_{\text{dice}}$ is the Dice loss [48, 56].
@@
STTran [8] ICCV’21 ResNet-101 25.2 34.1 37.0 68.6 71.8 71.8 24.6 36.2 48.8 77.9 94.2 99.1
APT [30] CVPR’22 ResNet-101 26.3 36.1 38.3 69.4 73.8 73.8 25.7 37.9 50.1 78.5 95.1 99.2
STTran-TPI [55] ACM MM’22 ResNet-101 26.2 34.6 37.4 69.7 72.6 72.6 - - - - - -
TR2 [53] ICRA’23 ResNet-101 26.8 35.5 38.3 70.9 73.8 73.8 27.8 39.2 50.0 83.1 96.6 99.9
VsCGG [35] ACM MM’23 ResNet-101 27.4 35.8 38.2 70.1 73.4 73.5 29.3 40.2 48.9 78.8 94.9 99.2
TEMPURA [36] CVPR’23 ResNet-101 28.1 33.4 34.9 68.8 71.5 71.5 29.8 38.1 46.4 80.4 94.2 99.4
DSG-DETR [13] WACV’23 ResNet-101 30.3 34.8 36.1 - - - 32.1 40.9 48.3 - - -
TPT [64] TMM’23 ResNet-101 - - - - - - 32.0 39.6 51.5 85.6 97.4 99.9
@@
OED [51] CVPR’24 ResNet-50 33.5 40.9 48.9 73.0 76.1 76.1 35.3 44.0 51.8 83.3 95.3 99.2
@@
Table 3: Performance comparison with SOTA PVSG methods on PVSG dataset. The best results are in bold. Next, $\diamondsuit$ and $\heartsuit$ stands for the relation predictor [62]: 1D Convolution and Transformer Encoder, respectively.
@@
Image Panoptic Segmentation + Tracking [4, 57]
@@
Video Panoptic Segmentation [4, 27]
@@ 4.1 Experimental Settings
Datasets & Evaluation Metrics. We conduct experiments on the Action Genome [19] for the DSGG task and the PVSG [62] for the PVSG task. We evaluate the DSGG task following the setting from [51] and the PVSG task following the setting from [62]. We train and evaluate the model strictly separately for each benchmark/task, without any data mixing or augmentation strategies.
Implementation Details. We adopt both Vision Transformer (ViT) and Convolutional Neural Networks (CNNs) as the backbone for the Visual Encoder to perform frame-wise feature extraction. More specifically, for ViT, we use ViT-S/14, ViT-B/14, ViT-L/14 [11] with pre-trained weights from DINO [39]; for CNNs, we use ResNet-50 [17] pre-trained weights from MoCo [16]. The number of slots are empirically chosen and can be observed that the optimal number of slots is influenced by both the number of object classes in the dataset and the objects present in the video. Therefore, for DSGG, we set $N=40$ object slots and $K=24$ relation slots; for PVSG, we use $N=96$ object slots and $K=40$ relation slots.
@@ 4.2 Comparison with State of the Arts
Results on Pixel-Level VidSGG (PVSG). The PVSG dataset presents highly dynamic videos and frequent substantial changes in camera angles. Tab. 3 demonstrates that both the Image Panoptic Segmentation + Tracking [4, 57] (IPS+T) model and Video Panoptic Segmentation [4, 27] (VPS) baselines fall short compared to our end-to-end UNO. It is essential to focus on R/mR@20, as it serves as a primary performance metric [62]. Notably, our model substantially outperforms both IPS+T and VPS at R/mR@20 across mask overlaps of 0.5 and 0.1. For instance, UNO holds the current peak for R@20 at 9.44% in the vIOU threshold of 0.5, indicating that, on average, 01 in every 13 ground-truth triplets is successfully recalled, compared to IPS+T’s best performance of 01 in roughly every 25 triplets (R@20 of 4.51%). However, lowering the threshold to a more lenient 0.1 raises UNO’s score to approximately 17.54%, allowing the model to recall 02 out of every 13 triplets. This suggests that while the model shows higher efficacy than others for recognizing key video content, there remains considerable room for improvement.
@@ 4.3 Ablation Studies
Effect of Multi-Task Learning. Tab. 4 highlights the impact of multi-task learning and cross-task synergy on PVSG, which provides both box-level and pixel-level VidSGG ground truths. AP50 denotes Average Precision@0.5, while PQ is the Panoptic Quality [22]. Using UNO, we observe that training with both bounding boxes and masks improves all metrics, with AP50 increasing from 28.5 to 29.6, PQ from 47.4 to 48.3, and R@20 from 8.03 to 9.44. These results suggest that box-level and pixel-level VidSGG data complement each other, enabling richer representations that enhance individual task performances.